Felix is purchasing a brownstone townhouse for ​$2,900,000. To obtain the​ mortgage, Felix is required to make a 19​% down payment. Felix obtains a 25​-year mortgage with an interest rate of 5.5​%.
​a) Determine the amount of the required down payment.
​b) Determine the amount of the mortgage.
​c) Determine the monthly payment for principal and interest.

Answer by Boreal(15235)

down payment is 2900000*0.19=$551,000
mortgage is 2900000-551000=2349000
monthly payment=mortgage(1+r/n)^300(r/n)/{[1+(r/n)]^300-1}, the 300 is number of months in 25 years
round at end
1+r/n=1.04583
This will yield $14424.92 a month monthly payment